For a SKF Single-Row 02-80 mm Deep-Groove ball bearing, an axial load Fa of 7650 N and a radial load Fr of 15 kN are applied with the

inner ring stationary. Estimate the bearing life at a speed of 740

rev/min with a reliability of 99.99%. SKF rates its bearings for 1 million

revolutions. The Weibull parameters are, x0 = 0.02, θ = 4.459, and b =

1.483


1
Expert's answer
2021-07-12T03:36:58-0400

"\frac{f_oF_a}{C_{or}}=\frac{13.2*1000}{24.0*10^3}=0.550\\
e=0.22+(0.26-0.22)*\frac{0.550-0.345}{0.689-0.345}=0.24 \\
\frac{F_a}{F_r}=\frac{1000}{3500}=0.29 >e\\
Resultant \\
X=0.56\\
Y=1.99-(1.99-1.71)* \frac{0.550-0.345}{0.689-0.345}=1.82 \\
L_{10h}=\frac{10^6}{60n}(\frac{C}{P})^p\\
L_{10h}=\frac{10^6}{60*800}(\frac{50.9*10^3}{3780})^3=50900 h\\"